# Extrapolation for bilinear compact operators in the variable exponent setting

### v1

@@ 5.A. Banach function spaces
We begin with briefly recalling the notion of Banach function spaces. We follow the presentation and definitions from the recent paper [25] of Lorist–Nieraeth, where we refer to for more details, history and proofs.
@@
Let $(\Omega,\mu)$ be a $\sigma$-finite measure space. We denote by $L^{0}(\Omega)$ the space of all measurable functions on $\Omega$. Let $X$ be a Banach space whose underlying set is a subset of $L^{0}(\Omega)$. Following [25], the space $X$ is said to be a Banach function space over $(\Omega,\mu)$ if it satisfies both of the following two properties:
@@
In several places of the literature, the definition of Banach function spaces includes a much stronger property than the saturation property, namely, that the characteristic function of any measurable set of finite measure belongs to the space. However, as explained in [25], this property is too restrictive, as it excludes in general for example weighted Lebesgue spaces, even with constant exponent.
Several important Banach functions spaces enjoy in addition the Fatou property. While sometimes in the literature it furnished part of the definition of Banach functions spaces, for us, following [25], it will be an additional, desirable property. We follow the formulation in [25].
@@
Another property that will be important for us is that of the absolutely continuous norm. We follow again the formulation in [25].
@@
First of all, it is well-known that $L^{p(\cdot)}(\mathbb{R}^{n})$ is a Banach function space over $\mathbb{R}^{n}$ that satisfies the Fatou property, see for instance [13, Lemma 3.2.8 (b)], and has absolutely continuous norm, see for example [12, p. 73]. Then, by [25, Proposition 3.17] we immediately obtain that $L^{p(\cdot)}(w)$ is a Banach function space over $\Omega$ that has the Fatou property. Furthermore, by combining [25, Proposition 3.12] with [25, Proposition 3.17] we deduce that $L^{p(\cdot)}(w)$ has absolutely continuous norm. Lastly, it is well-known that the Köthe dual of $L^{p(\cdot)}(\mathbb{R}^{n})$ (see [25] for the precise, general definition of the Köthe dual) is
@@
see for instance [12, p. 73]. Thus, by [25, Proposition 3.17] we obtain that the Köthe dual of $L^{p(\cdot)}(w)$ is
@@
In particular, $(L^{p(\cdot)}(w))^{\prime}$ has also absolutely continuous norm. Combining [25, Proposition 3.12] with [25, Corollary 3.16] we immediately deduce that $L^{p(\cdot)}(w)$ is reflexive. From [25, Proposition 3.15 (a)] we finally obtain that the Banach dual of $L^{p(\cdot)}(w)$ is $L^{p^{\prime}(\cdot)}(w^{-1})$. ∎
As proved in [25, Proposition 2.5], the saturation property has several equivalent reformulations. One of them is so important for us that we state it explicitly:
@@
By [25, Proposition 2.5 (ii)] we have that there exists an increasing sequence of measurable sets $(F_{k})^{\infty}_{k=1}$ such that $\chi_{F_{k}}\in X$ for all $k=1,2,\ldots$, and $\Omega=\bigcup_{k=1}^{\infty}F_{k}$. Moreover, since $\Omega$ is $\sigma$-finite, we can write $\Omega=\bigcup_{k=1}^{\infty}E_{k}$, where $(E_{k})^{\infty}_{k=1}$ is an increasing sequence of measurable sets with $\mu(E_{k})<\infty$, for all $k=1,2,\ldots$. Set $A_{k}:=F_{k}\cap E_{k}$, $k=1,2,\ldots$. Then, $(A_{k})^{\infty}_{k=1}$ is an increasing sequence of measurable sets with $\mu(A_{k})<\infty$, for all $k=1,2,\ldots$ and $\Omega=\bigcup_{k=1}^{\infty}A_{k}$. Moreover, since $0\leq\chi_{A_{k}}\leq\chi_{F_{k}}$, the ideal property yields $\chi_{A_{k}}\in X$, for all $k=1,2,\ldots$. ∎
@@
Let $(\Omega,\mu)$ be a $\sigma$-finite measure space. Following [20], we say that a sequence $(f_{k})^{\infty}_{k=1}$ of measurable functions on $\Omega$ converges locally in measure to a measurable function $f$ on $\Omega$, if for any measurable set $E\subseteq\Omega$ with $\mu(E)<\infty$ we have $f_{k}|_{E}\to f|_{E}$ as $k\to\infty$ in measure.
The local convergence in measure was characterized through a metric on the space of measurable functions in [20, Proposition B.6].
@@
By [20, Proposition B.6] we have that $d$ is a well-defined metric on $L^{0}(\Omega)$ making it into a topological vector space. Let $(f_{k})^{\infty}_{k=1}$ be a sequence of measurable functions on $\Omega$ and let $f\in L^{0}(\Omega)$. Then, the following are equivalent:
@@
That (a) implies (b) is just direction $\Rightarrow$ in [20, Proposition B.6 (iii)]. It is clear that (b) implies (c). Finally, the proof of direction $\Leftarrow$ in [20, Proposition B.6 (iii)] shows that (c) implies (a), since in that proof one only uses that $f_{m}|_{A_{k}}\to f|_{A_{k}}$ as $m\to\infty$ in measure, for all $k=1,2,\ldots$. ∎
@@
This is proved identically to [20, Proposition 3.3.1.3], by using Lemma 5.8 (c) in tandem with Lemma 5.9, instead of assumption [20, Definition 3.3.1.1 (iv)]. ∎
@@ 5.B. The Cobos–Fernández-Cabrera–Martínez theorem
We now recall the result of Cobos–Fernández-Cabrera–Martínez [7] about complex interpolation for bilinear compact operators. To state it, we need two additional pieces of notations: if $\bar{A}=(A_{0},A_{1})$ is a Banach couple, the we denote by $A_{j}^{\circ}$ the norm topological closure of $A_{0}\cap A_{1}$ in $A_{j}$, for all $j=0,1$. Moreover, we denote by ${\bf\mathcal{B}}(\bar{A}\times\bar{B},\bar{E})$ the family of bilinear operators $T$ defined on $(A_{0}\cap A_{1})\times(B_{0}\cap B_{1})$ with values in $E_{0}\cap E_{1}$ that satisfy the following
@@
In the original statement of [7, Theorem 3.2], the compactness property was assumed for $T:A_{0}^{\circ}\times B_{0}^{\circ}\to E_{0}$. However, because of the symmetry of the interpolation operator (see for example [3, Theorem 4.2.1]), this is equivalent to the formulation given here.
The definition of Banach function spaces in [7] includes more assumptions than our own. However, a careful examination of the very detailed exposition of the proof of [7, Theorem 3.2] in [20] (and the references therein) shows that the definition of Banach function spaces assumed here is sufficient.
In particular, let us remark that whenever one needs a sequence $(A_{k})^{\infty}_{k=1}$ of measurable sets having finite measure such that $\chi_{A_{k}}$ lies in the considered Banach space, for al $k=1,2,\ldots$ and $\bigcup_{k=1}^{\infty}A_{k}=\Omega$, such as in the proof of [20, Proposition 3.3.3.5], one can use Lemma 5.6. That is, one does not need to assume that the characteristic function of any measurable set of finite measure lies in the Banach space.
@@
6. Proof of abstract main result
In this section we finalize the proof of our abstract main result, Theorem 1.1.
The last ingredient we need is the following interpolation result which can be found in [30] (see also [15, Theorem 3.2]).
6.1 Theorem ([30], Theorems 3.4.1 and 3.4.2).
@@ 6. Proof of abstract main result
The statement of [30, Theorem 3.4.1] includes also the assumptions that $w_{j}\in L_{\text{loc}}^{p_{j}(\cdot)}$ and $w_{j}^{-1}\in L_{\text{loc}}^{p_{j}^{\prime}(\cdot)}$. However, in view of Lemma 5.5, a careful examination of the proof given there shows that this assumption is superfluous.
Also, the statement of [30, Theorem 3.4.2] includes the assumptions that $w_{j}\in L_{\text{loc}}^{p_{j}(\cdot)}$ and $w_{j}^{-1}\in L_{\text{loc}}^{p_{j}^{\prime}(\cdot)}$. However, as commented in the proof of [15, Theorem 3.2], the only property used in the proof is that weighted variable Lebesgue spaces are Banach lattices, which in view of Lemma 5.5 does not require these extra assumptions.
We have now all ingredients to prove Theorem 1.1.
@@
We prove the theorem in the case $t\in(1,\infty)$. The proof of the remaining case $t=1$ follows by similar arguments with the main difference that below one applies Lemma 4.1 instead of Lemma 4.2.
Let $\vec{p}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{2}$, $q(\cdot)\in\mathcal{E}^{0}_{(1,\infty)}$ such that
@@
Let also $\vec{w}$ be a 2-tuple of weights with $\vec{w}\in\mathcal{A}^{\gamma}_{\vec{p}(\cdot),q(\cdot)}$ and $\vec{w}^{t}\in\mathcal{A}^{t\gamma}_{\frac{\vec{p}(\cdot)}{t},\frac{q(\cdot)}{t}}$. We will show that $T:L^{p_{1}(\cdot)}(w_{1})\times L^{p_{2}(\cdot)}(w_{2})\to L^{q(\cdot)}(\nu_{\vec{w}})$ compactly.
To see that, we first apply Lemma 4.2: It delivers $\theta>0$, $\vec{p}_{0}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{2}$, $q_{0}(\cdot)\in\mathcal{E}^{0}_{(1,\infty)}$, such that
@@
$\frac{1}{\vec{p}(\cdot)}=\frac{1-\theta}{\vec{p}_{0}(\cdot)}+\frac{\theta}{\vec{p}_{1}(\cdot)},\quad\frac{1}{q(\cdot)}=\frac{1-\theta}{q_{0}(\cdot)}+\frac{\theta}{q_{1}(\cdot)},\quad\frac{1}{p_{0}(\cdot)}-\frac{1}{q_{0}(\cdot)}=\gamma,$
and a 2-tuple of weights $\vec{w}_{0}$ with $\vec{w}=\vec{w}_{0}^{1-\theta}\vec{w}_{1}^{\theta}$, such that $\vec{w}_{0}\in\mathcal{A}^{\gamma}_{\vec{p}_{0}(\cdot),q_{0}(\cdot)}$ and $\vec{w}_{0}^{t}\in\mathcal{A}^{t\gamma}_{\frac{\vec{p}_{0}(\cdot)}{t},\frac{q_{0}(\cdot)}{t}}$.
@@
By Lemma 5.5 and Remark 5.14 we have that $A_{j},B_{j},E_{j}$, $j=0,1$ satisfy all assumptions of Theorem 5.12. Moreover, since by assumption the operator $T:A_{1}\times B_{1}\to E_{1}$ is compact and $A_{1}^{\circ}$, respectively $B_{1}^{\circ}$, is a subspace of $A_{1}$, respectively $B_{1}$, we easily deduce from the equivalent characterizations of compactness of bilinear operators (see for instance [2, Proposition 1]) that also the restriction $T:A_{1}^{\circ}\times B_{1}^{\circ}\to E_{1}$ is compact. Therefore, Theorem 5.12 yields
@@ 7. Applications
In this section, we provide several applications of our compact extrapolation Theorem 1.1 for bilinear operators. In particular, we will show the compactness of commutators of bilinear $\omega$-Calderón–Zygmund operators and bilinear fractional integral operators.

### v2

@@ 5.A. Banach function spaces
We begin with briefly recalling the notion of Banach function spaces. We follow the presentation and definitions from the recent paper [undefz] of Lorist–Nieraeth, where we refer to for more details, history and proofs.
@@
Let $(\Omega,\mu)$ be a $\sigma$-finite measure space. We denote by $L^{0}(\Omega)$ the space of all measurable functions on $\Omega$. Let $X$ be a Banach space whose underlying set is a subset of $L^{0}(\Omega)$. Following [undefz], the space $X$ is said to be a Banach function space over $(\Omega,\mu)$ if it satisfies both of the following two properties:
@@
In several places of the literature, the definition of Banach function spaces includes a much stronger property than the saturation property, namely, that the characteristic function of any measurable set of finite measure belongs to the space. However, as explained in [undefz], this property is too restrictive, as it excludes in general for example weighted Lebesgue spaces, even with constant exponent.
Several important Banach functions spaces enjoy in addition the Fatou property. While sometimes in the literature it furnished part of the definition of Banach functions spaces, for us, following [undefz], it will be an additional, desirable property. We follow the formulation in [undefz].
@@
Another property that will be important for us is that of the absolutely continuous norm. We follow again the formulation in [undefz].
@@
First of all, it is well-known that $L^{p(\cdot)}(\mathbb{R}^{n})$ is a Banach function space over $\mathbb{R}^{n}$ that satisfies the Fatou property, see for instance [undefl, Lemma 3.2.8 (b)], and has absolutely continuous norm, see for example [undefk, p. 73]. Then, by [undefz, Proposition 3.17] we immediately obtain that $L^{p(\cdot)}(w)$ is a Banach function space over $\Omega$ that has the Fatou property. Furthermore, by combining [undefz, Proposition 3.12] with [undefz, Proposition 3.17] we deduce that $L^{p(\cdot)}(w)$ has absolutely continuous norm. Lastly, it is well-known that the Köthe dual of $L^{p(\cdot)}(\mathbb{R}^{n})$ (see [undefz] for the precise, general definition of the Köthe dual) is
@@
see for instance [undefk, p. 73]. Thus, by [undefz, Proposition 3.17] we obtain that the Köthe dual of $L^{p(\cdot)}(w)$ is
@@
In particular, $(L^{p(\cdot)}(w))^{\prime}$ has also absolutely continuous norm. Combining [undefz, Proposition 3.12] with [undefz, Corollary 3.16] we immediately deduce that $L^{p(\cdot)}(w)$ is reflexive. From [undefz, Proposition 3.15 (a)] we finally obtain that the Banach dual of $L^{p(\cdot)}(w)$ is $L^{p^{\prime}(\cdot)}(w^{-1})$. ∎
As proved in [undefz, Proposition 2.5], the saturation property has several equivalent reformulations. One of them is so important for us that we state it explicitly:
@@
By [undefz, Proposition 2.5 (ii)] we have that there exists an increasing sequence of measurable sets $(F_{k})^{\infty}_{k=1}$ such that $\chi_{F_{k}}\in X$ for all $k=1,2,\ldots$, and $\Omega=\bigcup_{k=1}^{\infty}F_{k}$. Moreover, since $\Omega$ is $\sigma$-finite, we can write $\Omega=\bigcup_{k=1}^{\infty}E_{k}$, where $(E_{k})^{\infty}_{k=1}$ is an increasing sequence of measurable sets with $\mu(E_{k})<\infty$, for all $k=1,2,\ldots$. Set $A_{k}:=F_{k}\cap E_{k}$, $k=1,2,\ldots$. Then, $(A_{k})^{\infty}_{k=1}$ is an increasing sequence of measurable sets with $\mu(A_{k})<\infty$, for all $k=1,2,\ldots$ and $\Omega=\bigcup_{k=1}^{\infty}A_{k}$. Moreover, since $0\leq\chi_{A_{k}}\leq\chi_{F_{k}}$, the ideal property yields $\chi_{A_{k}}\in X$, for all $k=1,2,\ldots$. ∎
@@
Let $(\Omega,\mu)$ be a $\sigma$-finite measure space. Following [undeft], we say that a sequence $(f_{k})^{\infty}_{k=1}$ of measurable functions on $\Omega$ converges locally in measure to a measurable function $f$ on $\Omega$, if for any measurable set $E\subseteq\Omega$ with $\mu(E)<\infty$ we have $f_{k}|_{E}\to f|_{E}$ as $k\to\infty$ in measure.
The local convergence in measure was characterized through a metric on the space of measurable functions in [undeft, Proposition B.6].
@@
By [undeft, Proposition B.6] we have that $d$ is a well-defined metric on $L^{0}(\Omega)$ making it into a topological vector space. Let $(f_{k})^{\infty}_{k=1}$ be a sequence of measurable functions on $\Omega$ and let $f\in L^{0}(\Omega)$. Then, the following are equivalent:
@@
That (a) implies (b) is just direction $\Rightarrow$ in [undeft, Proposition B.6 (iii)]. It is clear that (b) implies (c). Finally, the proof of direction $\Leftarrow$ in [undeft, Proposition B.6 (iii)] shows that (c) implies (a), since in that proof one only uses that $f_{m}|_{A_{k}}\to f|_{A_{k}}$ as $m\to\infty$ in measure, for all $k=1,2,\ldots$. ∎
@@
This is proved identically to [undeft, Proposition 3.3.1.3], by using Lemma 5.8 (c) in tandem with Lemma 5.9, instead of assumption [undeft, Definition 3.3.1.1 (iv)]. ∎
@@ 5.B. The Cobos–Fernández-Cabrera–Martínez theorem
We now recall the result of Cobos–Fernández-Cabrera–Martínez [undeff] about complex interpolation for bilinear compact operators. To state it, we need two additional pieces of notations: if $\bar{A}=(A_{0},A_{1})$ is a Banach couple, the we denote by $A_{j}^{\circ}$ the norm topological closure of $A_{0}\cap A_{1}$ in $A_{j}$, for all $j=0,1$. Moreover, we denote by ${\bf\mathcal{B}}(\bar{A}\times\bar{B},\bar{E})$ the family of bilinear operators $T$ defined on $(A_{0}\cap A_{1})\times(B_{0}\cap B_{1})$ with values in $E_{0}\cap E_{1}$ that satisfy the following
@@
In the original statement of [undeff, Theorem 3.2], the compactness property was assumed for $T:A_{0}^{\circ}\times B_{0}^{\circ}\to E_{0}$. However, because of the symmetry of the interpolation operator (see for example [undefb, Theorem 4.2.1]), this is equivalent to the formulation given here.
The definition of Banach function spaces in [undeff] includes more assumptions than our own. However, a careful examination of the very detailed exposition of the proof of [undeff, Theorem 3.2] in [undeft] (and the references therein) shows that the definition of Banach function spaces assumed here is sufficient.
In particular, let us remark that whenever one needs a sequence $(A_{k})^{\infty}_{k=1}$ of measurable sets having finite measure such that $\chi_{A_{k}}$ lies in the considered Banach space, for al $k=1,2,\ldots$ and $\bigcup_{k=1}^{\infty}A_{k}=\Omega$, such as in the proof of [undeft, Proposition 3.3.3.5], one can use Lemma 5.6. That is, one does not need to assume that the characteristic function of any measurable set of finite measure lies in the Banach space.
@@
6. Proof of abstract main results
In this section we finalize the proofs of our abstract main results, Theorems 1.1 and 1.2.
The last ingredient we need is the following interpolation result which can be found in [undefae] (see also [undefn, Theorem 3.2]).
6.1 Theorem ([undefae], Theorems 3.4.1 and 3.4.2).
@@ 6. Proof of abstract main results
The statement of [undefae, Theorem 3.4.1] includes also the assumptions that $w_{j}\in L_{\text{loc}}^{p_{j}(\cdot)}$ and $w_{j}^{-1}\in L_{\text{loc}}^{p_{j}^{\prime}(\cdot)}$. However, in view of Lemma 5.5, a careful examination of the proof given there shows that this assumption is superfluous.
Also, the statement of [undefae, Theorem 3.4.2] includes the assumptions that $w_{j}\in L_{\text{loc}}^{p_{j}(\cdot)}$ and $w_{j}^{-1}\in L_{\text{loc}}^{p_{j}^{\prime}(\cdot)}$. However, as commented in the proof of [undefn, Theorem 3.2], the only property used in the proof is that weighted variable Lebesgue spaces are Banach lattices, which in view of Lemma 5.5 does not require these extra assumptions.
We have now all ingredients to prove Theorems 1.1 and 1.2.
@@
Let $(\vec{p}(\cdot),q(\cdot),\vec{1},\infty)$ be a proper $2$-admissible quadruple such that
$q_{-}>1,$
@@
Let also $\vec{w}$ be a 2-tuple of weights with $\vec{w}\in\mathcal{A}_{\vec{p}(\cdot),q(\cdot)}$ and $\vec{w}^{t}\in\mathcal{A}_{\frac{\vec{p}(\cdot)}{t},\frac{q(\cdot)}{t}}$. We will show that $T:L^{p_{1}(\cdot)}(w_{1})\times L^{p_{2}(\cdot)}(w_{2})\to L^{q(\cdot)}(\nu_{\vec{w}})$ compactly.
To see that, we first apply Lemma 4.3: It delivers $\theta>0$ and a proper $2$-admissible quadruple $(\vec{p}_{0}(\cdot),q_{0}(\cdot),\vec{1},\infty)$, such that
$(q_{0})_{-}>1,$
@@
$\frac{1}{\vec{p}(\cdot)}=\frac{1-\theta}{\vec{p}_{0}(\cdot)}+\frac{\theta}{\vec{p}_{1}(\cdot)},\quad\frac{1}{q(\cdot)}=\frac{1-\theta}{q_{0}(\cdot)}+\frac{\theta}{q_{1}(\cdot)}$
and
$\quad\frac{1}{p_{0}(\cdot)}-\frac{1}{q_{0}(\cdot)}=\gamma,$
as well as a 2-tuple of weights $\vec{w}_{0}$ with
$\vec{w}=\vec{w}_{0}^{1-\theta}\vec{w}_{1}^{\theta}$
such that $\vec{w}_{0}\in\mathcal{A}_{\vec{p}_{0}(\cdot),q_{0}(\cdot)}$ and $\vec{w}_{0}^{t}\in\mathcal{A}_{\frac{\vec{p}_{0}(\cdot)}{t},\frac{q_{0}(\cdot)}{t}}$.
@@
By Lemma 5.5 and Remark 5.14 we have that $A_{j},B_{j},E_{j}$, $j=0,1$ satisfy all assumptions of Theorem 5.12. Moreover, since by assumption the operator $T:A_{1}\times B_{1}\to E_{1}$ is compact and $A_{1}^{\circ}$, respectively $B_{1}^{\circ}$, is a subspace of $A_{1}$, respectively $B_{1}$, we easily deduce from the equivalent characterizations of compactness of bilinear operators (see for instance [undefa, Proposition 1]) that also the restriction $T:A_{1}^{\circ}\times B_{1}^{\circ}\to E_{1}$ is compact. Therefore, Theorem 5.12 yields
@@
Proof of Theorem 1.2.
Let $(\vec{p}(\cdot),q(\cdot),\vec{r},s)$ be a proper $2$-admissible quadruple such that
$q_{-}>1$
and
$\frac{1}{p(\cdot)}-\frac{1}{q(\cdot)}=\gamma.$
Let also $\vec{w}$ be a 2-tuple of weights with $\vec{w}\in\mathcal{A}_{(\vec{p}(\cdot),q(\cdot),\vec{r},s)}$. We will show that $T:L^{p_{1}(\cdot)}(w_{1})\times L^{p_{2}(\cdot)}(w_{2})\to L^{q(\cdot)}(\nu_{\vec{w}})$ compactly.
To see that, we first apply Lemma 4.1: It delivers $\theta>0$ and a proper $2$-admissible quadruple $(\vec{p}_{0}(\cdot),q_{0}(\cdot),\vec{r},s)$, such that
$(q_{0})_{-}>1,$
$\frac{1}{\vec{p}(\cdot)}=\frac{1-\theta}{\vec{p}_{0}(\cdot)}+\frac{\theta}{\vec{p}_{1}(\cdot)},\quad\frac{1}{q(\cdot)}=\frac{1-\theta}{q_{0}(\cdot)}+\frac{\theta}{q_{1}(\cdot)}$
and
$\quad\frac{1}{p_{0}(\cdot)}-\frac{1}{q_{0}(\cdot)}=\gamma,$
as well as a 2-tuple of weights $\vec{w}_{0}$ with
$\vec{w}=\vec{w}_{0}^{1-\theta}\vec{w}_{1}^{\theta}$
such that $\vec{w}_{0}\in\mathcal{A}_{(\vec{p}_{0}(\cdot),q_{0}(\cdot),\vec{r},s)}$.
The rest of the proof is similar as the one of Theorem 1.1. ∎
@@ 7. Applications
In this section, we provide several applications of our compact extrapolation Theorems 1.1 and 1.2 for bilinear operators. In particular, we will show the compactness of commutators of bilinear $\omega$-Calderón–Zygmund operators, bilinear fractional integral operators and bilinear Fourier multipliers.